OpenStudy (anonymous):

Can you please check this question? What is the exact value of 7q^4 over sqrt(q^15) A) 7 sqrt (q) over q^4 B) 7 sqrt (q) over q^3 C) 7 sqrt (q) over q^6 D) 7 sqrt (q) over q^5 I think it's A.

OpenStudy (phi):

\[ \frac{7q^4}{\sqrt{q^{14} \cdot q}} = \frac{7q^4}{q^7 \cdot \sqrt{q}}= \frac{7}{q^3 \cdot \sqrt{q}}\] but people often "rationalize" a square root in the denominator. Multiply top and bottom by sqrt{q} \[ \frac{7\cdot \sqrt{q}}{q^3 \cdot \sqrt{q}\cdot \sqrt{q}}= \frac{7\cdot \sqrt{q}}{q^3 \cdot q}= \frac{7\cdot \sqrt{q}}{q^4}\]

OpenStudy (phi):

I think you are correct. It is A